OpenStudy (anonymous):

Solve x2 + 4x – 12 = 0 by completing the square. Show your work for full credit

OpenStudy (anonymous):

factorise: (x+6)(x-2)=0 therefore, either x+6=0 or x-2=0 x+6=0 x=-6 or x-2=0 x=2

OpenStudy (anonymous):

\[x^2 + 4x - 12 = 0\] \[(x + 2)^2 -12 - 4 =0\] \[(x + 2)^2 - 16 = 0\] \[(x+2)^2 = 16\] (x+2) = 4 OR (x+2)=-4 x = 2 OR x = -6

hero (hero):

Yeah I goofed up and forgot to add the 4 to both sides

OpenStudy (anonymous):

oh sorry, i didnt know what completing the square was :S